OpenStudy (aonz):

Halp please. If x=log2 and y=log3 express the following in terms of x and y.... log75.

OpenStudy (aonz):

@daniel987600

OpenStudy (raden):

log75 = log (25 * 3) = log 25 + log 3 = log(5^2) + log 3

OpenStudy (raden):

this is a start then use the property of logarithm : log a^m = m * log a so, log 5^2 = 2log 5 therefore log(5^2) + log 3 = 2 log 5 + log 3

OpenStudy (aonz):

yes i understand that but i dont undertstand how i write in terms of x and y

OpenStudy (raden):

then setting log 5 = log (10/2) use the property logarithm again : log (m/n) = log m - log n

OpenStudy (aonz):

OH YES!!! THANK YOU!!! :D :D :D

OpenStudy (aonz):

yesyes i understand it now :)

OpenStudy (raden):

so, log 5 = log (10/2) = log 10 - log 2 = 1 - log 2

OpenStudy (raden):

just subs the values log 2 = x and log 3 = y you're welcome :D